# A sparsity test for multivariate Hawkes processes

### v1

@@ 7.2. Proof of Theorem 1
First condition The limit function $\bar{\Lambda}$ is twice differentiable and satisfies
@@
Second condition For any $n$ in $\mathbb{N}^{\star}$ and $\theta$ in a neighbourhood of $\theta$,
(76) $\frac{1}{n}\boldsymbol{\Lambda}(n,\theta)=\boldsymbol{D}_{n}^{\textsc{t}}(\theta-\theta_{0})+\boldsymbol{R}_{n}(\theta)^{\textsc{t}}(\theta-\theta_{0}),$
where the remainder $\boldsymbol{R}_{n}(\theta)$ is stochastically equicontinuous in the sense that, denoting by $\bar{\boldsymbol{R}}$ the limit in probability under $\mathbb{P}(\theta_{0})$ of $n^{-1}\boldsymbol{R}(n,\cdot)$, for any $\varepsilon>0$ and $\eta>0$ we may find a neighbourhood $K$ of $\theta_{0}$ such that
(77) $\lim_{n\to\infty}\mathbb{P}\big{[}\sup_{\theta\in K}\sqrt{n}\big{\lVert}\boldsymbol{R}_{n}(\theta)-\boldsymbol{R}(\theta)\big{\rVert}>\eta\big{]}\leq\varepsilon.$
@@
where $\theta^{\star}$ lies on the line between $\theta_{0}$ and $\theta$ and is correctly defined on account of the convexity of $\Theta$. A sufficient condition for the stochastic equicontinuity of the remainder is therefore that $\partial_{\theta}\boldsymbol{\mathcal{S}}(n,\cdot)$ be $O_{\mathbb{P}(\theta_{0})}(1)$ uniformly in $\theta$. We use again the notation of Lemma 9, for all $s,\theta\in[0,T]\times\Theta$,
@@
$\partial_{\theta}\boldsymbol{\mathcal{S}}(n,\theta)=\Big{\{}\sum_{k=1}^{K}\int_{0}^{T}\Phi(Y_{k,ij}(s,\theta))\mathop{}\!\mathrm{d}N_{k}^{(n)}(s)-\int_{0}^{T}\partial_{\theta_{i}}\partial_{\theta_{j}}\lambda^{(n)}_{k,s}(\theta)\mathop{}\!\mathrm{d}s\Big{\}}_{1\leq i,j\leq d},$
@@
$\sqrt{n}(\frac{1}{n}\partial_{\theta}\boldsymbol{\mathcal{S}}(n,\theta)$ $-\partial_{\theta}\bar{\boldsymbol{\mathcal{S}}}(\theta,\theta_{0}))_{ij}=\sqrt{n}\sum_{k=1}^{K}\int_{0}^{T}\{\partial_{\theta_{i}}\partial_{\theta_{j}}\frac{1}{n}\lambda_{k,s}^{(n)}(\theta)-\partial_{\theta_{i}}\partial_{\theta_{j}}\bar{\lambda}_{k}(s,\theta,\theta_{0})\}\mathop{}\!\mathrm{d}s$
(79) $+\sqrt{n}\sum_{k=1}^{K}\int_{0}^{T}\Phi\big{(}\bar{Y}_{k,ij}(s,\theta,\theta_{0})\big{)}\{\frac{1}{n}\mathop{}\!\mathrm{d}N^{(n)}_{k}(s)-h_{k}(s,\theta,\theta_{0})\mathop{}\!\mathrm{d}s\}$
$+\sqrt{n}\sum_{k=1}^{K}\int_{0}^{T}\Phi\big{(}\frac{1}{n}Y_{k,ij}^{(n)}(s,\theta)\big{)}-\Phi\big{(}\bar{Y}_{k,ij}(s,\theta,\theta_{0})\big{)}\frac{1}{n}\mathop{}\!\mathrm{d}N^{(n)}_{k}(s).$
Consider the last term of (79).Re-arranging the terms again, for any $t\in[0,T]$,
$\sqrt{n}\Phi\big{(}\frac{1}{n}Y_{k,ij}^{(n)}(t,\theta)\big{)}$ $-\Phi\big{(}\bar{Y}_{k,ij}(t,\theta,\theta_{0})\big{)}=$
@@
$+$ $\frac{1}{\frac{1}{n}\lambda_{k}^{(n)}}\sqrt{n}\Big{\{}\frac{1}{n}\partial_{\theta_{i}}\partial_{\theta_{j}}\lambda_{k}^{(n)}(s,\theta)-\partial_{\theta_{i}}\partial_{\theta_{j}}\bar{\lambda}_{k}(s,\theta,\theta_{0})\Big{\}}$
@@
$\sum_{k=1}^{K}\frac{1}{\sqrt{n}}\int_{0}^{T}\Phi(\bar{Y}_{ij,k}(s,$ $\theta,\theta_{0}))\mathop{}\!\mathrm{d}M^{(n)}_{k}(s)$
(81) $+\int_{0}^{T}\Phi(\bar{Y}_{ij,k}(s,\theta,\theta_{0}))\sqrt{n}(\frac{1}{n}\lambda^{(n)}_{k,s}(\theta)-\bar{\lambda}_{k}(s,\theta,\theta_{0}))\mathop{}\!\mathrm{d}s.$
The last term of the sum (81) is uniformly $O_{\mathbb{P}}(1)$ as a direct consequence of Corollary 4. The deterministic function $\bar{Y}$ being differentiable by virtue of Assumption 6 and (44), we may once again integrate by parts as in the proof of Corollary 2. Hence the first term of (81) is uniformly bounded by
@@
(82) $\sqrt{n}\boldsymbol{D}_{n}\xrightarrow[n\to\infty]{\mathcal{L}(n,\theta_{0})}\mathcal{N}(0,A)$

### v2

@@ 7.2. Proof of Theorem 1
First condition. The limit function $\bar{\Lambda}$ is twice differentiable and satisfies
@@
Second condition. For any $n\in\mathbb{N}^{\star}$ and $\theta$ in a neighbourhood of $\theta$,
(76) $\Lambda_{n}(\theta_{0},\theta)=\boldsymbol{D}_{n}^{\textsc{t}}(\theta-\theta_{0})+\boldsymbol{R}_{n}(\theta)^{\textsc{t}}(\theta-\theta_{0}),$
where the remainder $\boldsymbol{R}_{n}(\theta)$ is stochastically equicontinuous in the sense that, denoting by $\bar{\boldsymbol{R}}$ the limit in probability under $\mathbb{P}(\theta_{0})$ of $n^{-1}\boldsymbol{R}_{n}$, for any $\varepsilon>0$ and $\eta>0$ we may find a neighbourhood $K$ of $\theta_{0}$ such that
(77) $\lim_{n\to\infty}\mathbb{P}\Big{[}\sup_{\theta\in K}\sqrt{n}\big{\lVert}\frac{1}{n}\boldsymbol{R}_{n}(\theta)-\boldsymbol{R}(\theta)\big{\rVert}>\eta\Big{]}\leq\varepsilon.$
@@
where $\theta^{\star}$ lies on the line between $\theta_{0}$ and $\theta$ and is correctly defined on account of the convexity of $\Theta$. A sufficient condition for the stochastic equicontinuity of the remainder is therefore that $\sqrt{n}\{n^{-1}\partial_{\theta}\boldsymbol{\mathcal{S}}(n,\theta)-\partial_{\theta}\bar{\boldsymbol{S}}(\theta,\theta_{0})\}$ be $O_{\mathbb{P}(\theta_{0})}(1)$ uniformly in $\theta$. We use again the notation of Lemma 9, for all $s,\theta\in[0,T]\times\Theta$,
@@
$\partial_{\theta}\boldsymbol{\mathcal{S}}(n,\theta)=\Big{\{}\sum_{k=1}^{K}\int_{0}^{T}\Phi\big{(}\frac{1}{n}Y^{(n)}_{k,ij}(s,\theta)\big{)}\mathop{}\!\mathrm{d}N_{k,s}^{(n)}-\int_{0}^{T}\partial_{\theta_{i}}\partial_{\theta_{j}}\lambda^{(n)}_{k,s}(\theta)\mathop{}\!\mathrm{d}s\Big{\}}_{1\leq i,j\leq d},$
@@
$\sqrt{n}(\frac{1}{n}\partial_{\theta}\boldsymbol{\mathcal{S}}(n,\theta)$ $-\partial_{\theta}\bar{\boldsymbol{\mathcal{S}}}(\theta,\theta_{0}))_{ij}=\sqrt{n}\sum_{k=1}^{K}\int_{0}^{T}\{\partial_{\theta_{i}}\partial_{\theta_{j}}\bar{\lambda}_{k}(s,\theta,\theta_{0})-\partial_{\theta_{i}}\partial_{\theta_{j}}\frac{1}{n}\lambda_{k,s}^{(n)}(\theta)\}\mathop{}\!\mathrm{d}s$
(79) $+\sqrt{n}\sum_{k=1}^{K}\int_{0}^{T}\Phi\big{(}\bar{Y}_{k,ij}(s,\theta,\theta_{0})\big{)}\{\frac{1}{n}\mathop{}\!\mathrm{d}N^{(n)}_{k,s}-h_{k}(s,\theta_{0})\mathop{}\!\mathrm{d}s\}$
$+\sqrt{n}\sum_{k=1}^{K}\int_{0}^{T}\Phi\big{(}\frac{1}{n}Y_{k,ij}^{(n)}(s,\theta)\big{)}-\Phi\big{(}\bar{Y}_{k,ij}(s,\theta,\theta_{0})\big{)}\frac{1}{n}\mathop{}\!\mathrm{d}N^{(n)}_{k,s}.$
Consider the last term of (79).Re-arranging the terms again, for any $s\in[0,T]$,
$\sqrt{n}\big{\{}\Phi\big{(}\frac{1}{n}Y_{k,ij}^{(n)}(s,\theta)\big{)}$ $-\Phi\big{(}\bar{Y}_{k,ij}(s,\theta,\theta_{0})\big{)}\big{\}}=$
@@
$+$ $\frac{1}{\frac{1}{n}\lambda_{k,s}^{(n)}(\theta)}\sqrt{n}\Big{\{}\frac{1}{n}\partial_{\theta_{i}}\partial_{\theta_{j}}\lambda_{k}^{(n)}(s,\theta)-\partial_{\theta_{i}}\partial_{\theta_{j}}\bar{\lambda}_{k}(s,\theta,\theta_{0})\Big{\}}$
@@
$\sum_{k=1}^{K}\frac{1}{\sqrt{n}}\int_{0}^{T}\Phi(\bar{Y}_{ij,k}(s,$ $\theta,\theta_{0}))\mathop{}\!\mathrm{d}M^{(n)}_{k,s}$
(81) $+\int_{0}^{T}\Phi(\bar{Y}_{ij,k}(s,\theta,\theta_{0}))\sqrt{n}(\frac{1}{n}\lambda^{(n)}_{k,s}(\theta_{0})-\bar{\lambda}_{k}(s,\theta_{0},\theta_{0}))\mathop{}\!\mathrm{d}s.$
The last term of the sum (81) is uniformly $O_{\mathbb{P}}(1)$ as a direct consequence of Corollary 4. The deterministic function $\bar{Y}$ being differentiable in $t$ by virtue of Assumption 6 and (44), we may once again integrate by parts as in the proof of Corollary 2. Hence the first term of (81) is uniformly bounded by
@@
(82) $\frac{1}{\sqrt{n}}\boldsymbol{D}_{n}\xrightarrow[n\to\infty]{\mathcal{L}(\mathbb{P}(\theta_{0}))}\mathcal{N}(0,A)$